Consider an electric circuit with an inductance of 5 henries

and a resistance of 12 ohms in series with an e.m.f 120sin200t

volts. Find the current


1
Expert's answer
2022-03-24T19:29:51-0400

Due to Kirchhoff’s law:

"L\frac {di}{dt}+Ri=e"

"5\frac {di}{dt}+12i=120\sin{200t}"

The solutions to a nonhomogeneous equation are of the form

"i(t) = i_c(t) + i_p(t)"

where "i_c(t)" is the general solution to the associated homogeneous equation and "i_p(t)" is a particular solution.

The associated homogeneous equation:

"5\frac {di}{dt}+12i=0"

"i_c(t)=Ce^{-\frac {12}{5}t}"

"i_p(t)=A\sin{200t}+B\cos{200t}"

"i_p’(t)=A\cdot200\cos{200t}-B\cdot200\sin{200t}"

Now put these into the original differential equation to get:

"5 (A\cdot200\cos{200t}-B\cdot200\sin{200t} )+" "12(A\sin{200t}+B\cos{200t})=120\sin{200t}"

"A=\frac{90}{62509}" ; "B=-\frac{7500}{62509}".

Answer: "i( t)=Ce^{-\frac {12}{5}t}+ \frac{90}{62509}\sin{200t}-\frac{7500}{62509}\cos{200t}"